Fermi National Laboratory


Accelerator Update for January 5, 2005 - January 7, 2005

Wednesday January 5
The day shift began with the Antiproton Source (Pbar) stacking, and with the following machines conducting studies: Booster, Recycler, Switchyard (SY), Tevatron (TeV), and MiniBooNE.

Main Injector (MI) RF technicians reported at 9 AM that MIRF13 was up and running, which means MI is back to 18 RF stations. Later, techs said that this station needed its modulator changed. This was done at 1 PM.

The Run Coordinator determined that stacking should be held off to repair a LCW leak on a Lambertson (LAM42) that drenched a transformer. Later, power supply techs said that this supply would have to dry out for days before it could be used again. At 3:20 PM, a new power supply was being moved over as a replacement.

At 8:33 PM, a D-Zero (D0) experimenter reported that they had a problem with their Muon system and need an hour and a half access to make repairs.

Operations established a 14x0 study-store at 8:49 PM

At 9:17 PM, a MI power supply (MBEX) tripped off. A monitor was set up to show what was tripping and operators went to the MI-10 Kicker room to read it. The power supply tripped off many times during the evening shift.

LAM42 work was suspended for the night; it will be completed on Thursday morning. Pbar began single batch stacking cycles at 10 PM.

Operators prepared the Collider Detector at Fermilab (CDF) and D0 for access at 11:16 PM.

Thursday January 6
The midnight shift began with Pbar stacking, with CDF and D0 in access, and with MiniBooNE taking beam.

D0 came out of access at 12:06 AM.

CDF came out of access at 12:26 AM.

Operations put the TeV through a dry squeeze at 12:54 AM.

Operations put the TeV into shot setup at 1:22 AM.

MBEX tripped again at 1:42 AM. It tripped off several more times during the midnight shift. Operations lowered intensity so that MiniBooNE could keep getting beam.

Operations established store 3906 at 4:39 AM with an initial luminosity of 54.06E30. Pbar resumed stacking.

At 4:45 AM, MBEX tripped off and wouldn't reset. Two experts were called in to make repairs.

Operations established beam to the Meson experiments MTest and MCenter at 4:57 AM.

Experts had the MBEX power supply up and running by 6:25 AM.

At 8:25 AM, the D0 solenoid was ramped down to work on a cryo system compressor.

Operations prepared the Recycler for stashing antiprotons at 9:25 AM.

Techs reported that LAM42 was up and running at 9:38 AM.

Operations reported three transfers of antiprotons to the Recycler.

Operations halted Pbar stacking at 10:36 AM to set up for the normal mode of Pbar slipstacking. Pbar resumed stacking at 10:44 AM.

At 7 PM, CDF reported a problem with the colliding beam. Operations contacted the TeV coordinator and operators made some changes.

Friday January 7
The midnight shift began with Operations monitoring store 3906, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

MTest finished using beam at 12:07 AM.

TeV expert began an end-of-store study at 6:18 AM.

Operations terminated store 3906 at 7:18 AM.

The Plan for Friday and the Weekend
The plan for today is to established a mixed mode Recycler and Pbar store. The plan for this weekend and from now on is to always create mixed Pbar mode stores.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/7/2005   email Fermilab